Classify the following signal as energy signal or power signal. Find the
normalized energy or normalized power.
(i) sin 2t + 3 cos 4t
1
Expert's answer
2021-08-18T09:29:01-0400

Checking whether the signal is energy or power

"\int _{-\infty \:}^{\infty \:}|x(t)|^2\\
\int _{-\infty \:}^{\infty \:}\left(\sin \left(2t\right)+3\cos \left(4t\right)\right)^2dt\\
=5t-\frac{1}{8}\sin \left(4t\right)-\frac{1}{2}\cos \left(6t\right)+\frac{3}{2}\cos \left(2t\right)+\frac{9}{16}\sin \left(8t\right)+C\\
=\infin"

The energy of the signal is infinite hence the signal is power.